OpenStudy (anonymous):

Need to find the slope-intercept equation of the line that has given characteristics and write in slope-intercept form slope -6 y-intercept (0,2)

OpenStudy (anonymous):

Slope-Intercept Form- Y=mx+b Simply plug in the slope and y-intercept. Slope- m Y-Intercept- b So your answer would be y=-6x+2

OpenStudy (anonymous):

y-2=-6(x-0) y=-6x+2